Netflix is selling for $100 a share. A Netflix call option with one month until expiration and an exercise price of $111 sells for $2.40 while a put with the same strike and expiration sells for $12.18. a. What is the market price of a zero-coupon bond with face value $111 and 1-month maturity?

a. What is the market price of a zero-coupon bond with face value $111 and 1-month maturity

b. What is the risk-free interest rate expressed as an effective annual yield?
Answer #1

Using put call parity
-C+P+S=X/(1+r)^t

1.
=X/(1+r)^t
=-2.40+12.18+100
=109.78

2.
109.78=111/(1+r)^(1/12)
=>r=(111/109.78)^12-1
=>r=14.182%